OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (anonymous):

@ganeshie8 @Directrix

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@ganeshie8

OpenStudy (agent0smith):

Find the magnitude. Divide the x and y components by said magnitude. That's your unit vector.

OpenStudy (johnweldon1993):

Indeed^ \[\large u = <u_x, u_y>\] \[\large ||u|| = \sqrt{(u_x)^2 + (u_y)^2}\] \[\large \hat{u} = \frac{<u_x,u_y>}{||u||} = <\frac{u_x}{||u||}, \frac{u_y}{||u||}>\]

OpenStudy (shamim):

U know u is a vector

OpenStudy (shamim):

X component of u vector is =-4

OpenStudy (shamim):

Y component of u vector is=-3

OpenStudy (shamim):

i is unit vector along x axis

OpenStudy (shamim):

j is unit vector along y axis

OpenStudy (shamim):

So vector u=-4i-3j Right?

OpenStudy (shamim):

Can u tell me the unit vector along vector u=?